ii) Suppose that a customer deposits $7,000 into their bank account. Based on the simple money multiplier calculated in part i), calculate the final change in the money supply available in the banking system. (Part i answer is 6.25)


iii) Calculate the final change in the money supply available in the banking system if the reserve ratio increases to 25%. Explain briefly how the money multiplier works.
1
Expert's answer
2018-01-09T05:04:53-0500
ii) If a customer deposits $7,000 into their bank account and simple money multiplier is m = 6.25, then the final change in the money supply available in the banking system is 7,000/rr = 7,000/(1/m) = 7,000/(1/6.25) = $43,750 (where m = 1/rr).

iii) If rr = 25%, then the final change in the money supply available in the banking system is $7,000/0.25 = $28,000.